# Mon CHÈRI $\heartsuit$ Adapting Capability Hardware Enhanced RISC with Conditional Capabilities

### v1

@@ IV-D Write-before-Read Support for Memory Allocators
To benefit from Write-before-Read for heap allocations, the dynamic heap allocator must be augmented to initialize the operation bounds of allocated memory areas with the appropriate CP. For implementing low-level support for conditional capabilities into system software like memory allocators, we implemented the cheri_bounds_set() application programming interface (API) that can be used by system programmers to interface directly with the Mon CHÈRI instructions. LABEL:lst:heap shows how we modified the CheriFreeRTOS memory allocator, and the TLSF allocator [15], previously ported to CHERI [63] to enforce Write-before-Read on allocation made by malloc(). Since these allocators had already been augmented to be CHERI-aware, the malloc() implementation already called the cheri_bounds_set intrinsic ➀ supported by CHERI-LLVM to set the bounds of the return pointer ptr to match the allocation size. To make them conditional capability-aware, we added a call to cheri_opbounds_set() intrinsic ➁ before malloc() returns. This sets the operation bounds of ptr explicitly to zero, and unsets the PERM_WRITE_BEFORE_READ_DISABLE control bit.
V Evalution
TABLE III: Detection rate of Mon CHÈRI on uninitialized memory issues from Juliet Test Suite [56] CWE457 test cases. Green cells indicate the true positives and true negatives, while red cells indicate the false positives and false negatives.
Ground truth Mon CHÈRI
@@ V Evalution
Following the “Good” and “Bad” classification in Juliet, Mon CHÈRI reports six false-positives. Yet, these cases do exhibit uninitialized memory access behavior (cf. § V-A).
@@
Write-before-Read + purecap 50224 6496 $14.85\%$ 2951842261 247562975 $9.15\%$ 29 2 $7.04\%$ 344 26 $7.00\%$ 3.44
@@
We evaluated conditional capabilities using our Mon CHÈRI prototypes for functionality, security, performance, and area cost. The functional and security evaluation, described in § V-A, was conducted on a conditional capability-enhanced QEMU-system-CHERI128 full-system emulator. We used CheriFreeRTOS [1], an augmented version of FreeRTOS with CHERI-based compartmentalization, incorporating a Write-before-Read memory allocator as detailed in § IV-D.
@@
For performance and area cost evaluation, we extended the CHERI-RISC-V field-programmable gate array (FPGA) softcore [18] based on the open-source Bluespec [54] RISC-V processor IP. This processor family includes Piccolo, Flute, and Tooba IP cores. Although sharing significant portions of Bluespec System Verilog (BSV) code, our performance evaluation uses the 64-bit CHERI-RISC-V Flute (RV64ACDFIMSUxCHERI) where we implemented the Mon CHÈRI Write-before-Read ISA extension. We refer to this prototype as MonCHÈRI-Flute64. We used the RISC-V ISA tests [53] to confirm the operation of MonCHÈRI-Flute64 is correct w.r.t RISC-V specifications. It passed all 229 tests.
V-A Functional and Security Evaluations on QEMU
The Juliet Test Suite [56] is a systematic set of thousands of small test programs in C/C++ that demonstrate common programming practices that lead to memory vulnerabilities. These are organized by Common Weakness Enumeration (CVE) numbers. Each sample comes in two versions: a “Bad” case that exhibits the vulnerability and “Good” case that has been patched. We evaluated Mon CHÈRI using a subset of CWE457 [51] (Use of Uninitialized Variable) from the Juliet Test Suite, which covers 560 bad and 560 corresponding good cases.
Juliet is designed primarily as a benchmark for static program analysis tools. A challenge with using the Juliet test cases as-is is that modern compilers will automatically reject a significant portions of the test cases thanks to static program analysis for uninitialized variables. Even when suppressing uninitialized-variable warnings, LLVM will exploit the undefined behavior of variables with uninitialized values to optimize away memory accesses. To ensure our tests exercise Mon CHÈRI on uninitialized memory accesses, we opted to suppress such compiler optimizations by declaring all variables as volatile. This allows us to build the tests with optimization level -O2 while ensuring all test cases lead to memory accesses. With these adaptations, we ported Juliet to operate within CheriFreeRTOS and to use CheriFreeRTOS’ instrumented allocation API. CheriFreeRTOS compartmentalization offers the “compare and return” mode. This functionality enables a compartment to detect and report a memory vulnerability and subsequently return control to the caller, enhancing error handling: we run every test case in a different compartment and catch the result.
The results, illustrated in Table III, show that Mon CHÈRI detects all bad cases (true-positive rate 100%) and miss-detects only six good cases (false-positive rate 1%). However, when looking into the reported false positives in detail, we discover that these test cases actually do contain uninitialized memory reads. In § -B, LABEL:lst:juliet-cwe457-63-good we provide the source code of one of these cases: A pointer to an uninitialized variable is passed as an argument to a function call; the callee then dereferences and copies the memory content of the variable, which Mon CHÈRI correctly detects as a violation of the Write-before-Read policy. The read memory content is not actually used in the test case but gets immediately overwritten (i.e., initialized). As is, the test case is actually safe, and analysis tools that rely, e.g., on taint propagation [65], may detect that the uninitialized memory content is neither propagated nor processed in any way. As Mon CHÈRI enforces memory access policies per instruction at the architectural level, we currently cannot distinguish this benign case of an uninitialized read from potential vulnerabilities. All six false positives reported in Table III follow the same pattern where uninitialized memory content is copied but not processed.
V-B Performance and Area Cost Evaluation on FPGA
@@
We synthesized MonCHÈRI-Flute64 at 100 MHz on an AMD Virtex UltraScale+ VCU-118 FPGA. Table IV shows the area cost of MonCHÈRI-Flute64 relative to the CHERI-Flute64 design. It has a $2\%$ increase in area size, which is low considering the cost of adding CHERI to the Flute IP [1].
Although MonCHÈRI-Flute64 only implements the Write-before-Read ISA extension, most of the additional logic affecting the processor pipeline (see Figure 3(b) in § III-B) is common for enforcing different CPs listed in Table I). In designs incorporating multiple CPs, the additional cost to the conditional capability ISA extension is expected to be mostly amortized by our implementation of Write-before-Read. A limiting factor is the limited unused bits in the CHERI capability representation (see Figure 4(a) in § IV-B). In § VII discuss potential solutions to this limitation in CHERI.
@@
For performance measurements, we integrated the MonCHÈRI-Flute64 softcore to the BESSPIN-GFE security evaluation platform[59]. The GFE system includes the MonCHÈRI-Flute64 softcore, a BootROM, Soft Reset and JTAG, UART, Ethernet/DMA, DDR4, and Flash controllers. A host-based gdb debugger connects to the system over the USB/JTAG connector, and a host-based console connects over USB/UART. For performance measurements, we ran the EEMBC CoreMark [22] benchmark bare metal on the GFE.
@@
Table V shows the performance results for MonCHÈRI-Flute64 relative to the CHERI-Flute64 design. We compared the results CHERI enforcement completely disabled (nocap), CHERI-enforcement in pure-capability mode (purecap), and Write-before-Read in MonCHÈRI-Flute64 enabled for all variables in CoreMark in pure-capability mode (Write-before-Read + purecap ❶). Our results show that Write-before-Read conditional capabilities add a small overhead compared to CHERI pure capability mode ($\approx$ 3.5%), while having a negligible impact on the baseline processor performance when turned off ($\approx$ 0.4%).
We also instrumented CoreMark without performing store linearization ❷. Although this configuration is not guaranteed to run, as explained in § IV-C, we disabled the hardware trap for failed operation bound checks to compare the performance impact of our hardware changes with that of the store linearization program transformation. We found that most of the performance degradation in the earlier experiments was due to program transformation, with the hardware changes having a negligible impact ($\approx$ 0.2%). We conclude that further performance improvements to Mon CHÈRI can be achieved by optimizing the store linearization strategy.
@@
Finally, we measure the performance impact on the TLSF allocator on MonCHÈRI-Flute64 through microbenchmarks allocating and freeing 1 MB of memory in chunks ranging from 32 bytes to 4 KiB. We compared TLSF performance in nocap, purecap, and Write-before-Read + purecap configurations. Additionally, we estimate the cost of zero-initializing allocated memory, similar to allocators performing automatic initialization. The results on Figure 5 show the Write-before-Read-enhanced TLSF allocator having negligible impact on performance: $\approx$ 0.1% (g.m.) compared to purecap and $\approx$ 1.5% (g.m.) compared to nocap. The overhead of Write-before-Read + purecap remains constant regardless of allocation size, whereas the overhead of zero-initialization scales linearly.
@@
Figure 5: TLSF allocator microbenchmark results. Bars show overhead relative to TLSF in MonCHÈRI-Flute64 nocap mode.
VI Related Work
@@
   Uninitialized capabilities [27] *3 ?
   Capstone [87] *3 ? $\approx\leavevmode\nobreak\ 50\%$4
   Mon CHÈRI $\approx$ 3.5% /\ /\ / 7%5

### v2

@@ 4.4 Mon CHÉRI Support for Memory Allocators
We implemented the cheri_opbounds_set() application programming interface (API) to provide low-level support for conditional capabilities in system software, such as memory allocators. To enable Write-before-Read CPs for heap allocations, the allocator must initialize the capabilities to allocate memory with the appropriate CP.
LABEL:lst:heap shows how we modified the CheriFreeRTOS memory allocator, and the TLSF allocator [32], previously ported to CHERI [42] to enforce Write-before-Read on allocation made by malloc(). A CHERI-aware malloc() already uses the cheri_bounds_set() intrinsic to set bounds of the return pointer, ptr, based on the allocation size. To make it conditional capability-aware, we added a call to cheri_opbounds_set(ptr, 0, WriteBeforeRead) before returning from malloc(). This explicitly sets the operation bounds of ptr to zero and configures the CP control bits to indicate ptr should be treated as Write-before-Read by the hardware.
5 Evaluation
TABLE III: Detection rate of Mon CHÉRI on uninitialized memory issues from Juliet Test Suite [43] CWE457 test cases. Green cells indicate the true positives and true negatives, while red cells indicate the false positives and false negatives.
Ground truth Mon CHÉRI
@@ 5 Evaluation
Following the “Good” and “Bad” classification in Juliet, Mon CHÉRI reports six false-positives. Yet, these cases do exhibit uninitialized memory access behavior (cf. § 5.1).
@@
Write-before-Read + purecap 50224 6496 $14.85\%$ 2949321554 245019752 $9.06\%$ 29 2 $7.04\%$ 344 26 $7.00\%$ 3.44
@@
We evaluated the Mon CHÉRI prototypes for functionality, security, performance, and area cost. The functional and security evaluation (§ 5.1) was performed on a conditional capability-enhanced QEMU-system-CHERI128 full-system emulator, CheriFreeRTOS [31], which integrates CHERI-based compartmentalization, and a Write-before-Read memory allocator (§ 4.4). For performance and area cost, we extended the CHERI-RISC-V field-programmable gate array (FPGA) softcore [44] based on the open-source Bluespec [45] RISC-V processor IP with the CP ISA extension (§ 4.1). This processor family includes Piccolo, Flute, and Tooba IP cores. Although sharing significant portions of Bluespec System Verilog (BSV) code, our performance evaluation uses the 64-bit CHERI-RISC-V Flute (RV64ACDFIMSUxCHERI). This prototype, MonCHÉRI-Flute64, was validated against RISC-V specifications using 229 RISC-V ISA tests [46].
5.1 Functional and Security Evaluation on QEMU
We used the U.S. National Institute of Standards and Technology (NIST) Juliet Test Suite [43], which includes thousands C/C++ of test cases that demonstrate common programming defects that lead to memory vulnerabilities, to evaluate Mon CHÉRI. These tests are organized by CVE numbers, with “bad” versions exhibiting vulnerabilities and “good” versions showing patched code. We focused on the CWE-457 [47] (Use of Uninitialized Variable) C test cases, covering 560 bad and 560 corresponding good cases. These examples cover a range of realistic scenarios, such as conditional control flows where variables might remain uninitialized in certain branches (e.g., as illustrated in LABEL:lst:example), function calls that pass variables assumed to be initialized, and cases involving complex data types like arrays, pointers, and structures.
To extensively assess Mon CHERI’s detection rate for uninitialized variable accesses, we made specific modifications to the Juliet test suite. Modern compilers like LLVM tend to optimize away uninitialized memory accesses or reject them outright due to their sophisticated static analysis capabilities. To prevent this behavior, we declared all variables as volatile, which stops the compiler from applying optimizations based on undefined behavior. This allows us to build the tests with optimization level -O2 while ensuring all test bad cases trigger uninitialized memory accesses.
The tests were executed within CheriFreeRTOS, which uses an instrumented allocation API to track memory accesses. CheriFreeRTOS provides a “compartmentalize and return” mode that isolates each test case in a separate compartment, allowing test cases resulting in a CHERI protection fault to return control to the caller, which records the result and proceeds to the next test case. The test cases were compiled with the Mon CHÉRI-enhanced CHERI-LLVM at optimization level -O2 and with Write-before-Read instrumentation and store linearization but without the Write-before-Read optimization described in § 4.3.
The results in Table III show that Mon CHÉRI detects all “bad” cases (100% true-positive rate) and reports only six false positives (1% false-positive rate). Upon closer inspection, these false positives stem from cases where uninitialized variables are copied, but never used. Although these cases are technically valid violations of the Write-before-Read policy, they do not pose a security risk, as data is immediately overwritten. Analysis tools relying on taint propagation [48] might not flag these cases, as the uninitialized memory does not propagate. As Mon CHÉRI enforces policies at an architectural level, distinguishing between these benign cases and actual vulnerabilities is not currently possible. All six false positives share this pattern, where the uninitialized memory is copied but later overwritten before being used. We provide the source code for one of these cases in § A.3, LABEL:lst:juliet-cwe457-63-good. Further, all six cases have straightforward software workarounds that can be applied, once detected, to initialize variables early with a default zero value to avoid uninitialized access.
We evaluated the effectiveness of the store linearization pass by comparing the detection rate for the Juliet tests instrumented with and without store linearization. Without store linearization 119 out of 560 “good” test cases exhibit false positives (21% false positive rate). This demonstrates store linearization provides a significant improvement to the detection accuracy of Mon CHÉRI. We also verified the Write-before-Read optimization did not affect the detection accuracy as it only omits CP instrumentation for variables that are statically verified to be fully initialized.
For comparison, we compiled the Juliet test suite with all relevant warnings enabled in GCC and Clang/LLVM. GCC detected 170 out of 560 uninitialized cases (30%) at -O0 and 173 cases (31%) at -O2. Clang/LLVM detected 117 cases (21%) regardless of optimization level. Detection rates for the Valgrind and Dr. Memory dynamic analysis tools range from below 10% to levels comparable to Mon CHÉRI, depending on the Juliet and compiler configurations. Under the same configuration used for Mon CHÉRI, (volatile variables and Clang with -O2), Valgrind and Dr. Memory detected 400 cases (71%) and 388 cases (69%), respectively.
As the Juliet CWE457 tests do not exhibit patterns that would require Write-before-Execute, Write-before-Read-Only, Write-before-Execute-Only, or Write-Once CPs we verified the functionality of these CPs in the conditional capability-enhanced QEMU-system-CHERI128 implementation using purpose-built synthetic test cases.
5.2 Performance and Area Evaluation on FPGA
@@
We synthesized MonCHÉRI-Flute64 at 100 MHz on an AMD Virtex UltraScale+ VCU-118 FPGA. Compared to the CHERI-Flute64 design, the area cost of MonCHÉRI-Flute64 increased by only 2%, a small cost considering the overhead of adding CHERI. The majority of this additional logic is shared across many CPs.
@@
We integrated the MonCHÉRI-Flute64 softcore into the BESSPIN-GFE security evaluation platform[49], which allows for a full-system evaluation of Mon CHÉRI performance. The GFE system includes the MonCHÉRI-Flute64 softcore, a BootROM, Soft Reset and JTAG, UART, Ethernet/DMA, DDR4, and Flash controllers. A host-based gdb debugger connects to the system over the USB/JTAG connector, and a host-based console connects over USB/UART. We measured performance using the EEMBC CoreMark [50] benchmark, running bare-metal on the GFE. Although MonCHÉRI-Flute64 supports all CPs in Table II we focus in these experiments on Write-before-Read as it is applicable to all variables in CoreMark. Consequently, applying it to all variables in the CoreMark benchmark code provides a worst-case estimate of Mon CHÉRI’s performance impact. As we expect other CPs to only be applied to a subset of variables, their impact is a fraction of Write-before-Read’s.
@@
Table V compares the performance results MonCHÉRI-Flute64 to the CHERI-Flute64 across different configurations: no capability enforcement (no-cap), pure-capability mode (purecap), and Write-before-Read enabled in pure-capability mode (Write-before-Read + purecap ❶). The results show that the Write-before-Read extension adds a modest $\approx$ 3.5% overhead over pure-capability mode, with minimal impact on baseline performance ($\approx$ 0.4%) when capability enforcement is disabled. The combined performance impact of Write-before-Read and CHERI pure-capability mode over the baseline performance with no capability enforcement is 7%.
We also compared the performance of MonCHÉRI-Flute64 with and without store linearization (❷). Although store linearization is necessary, as explained § 5.1, for the correctness of Write-before-Read enforcement, disabling the hardware fault for this experiment allows us to compare the performance impact of the hardware changes with that of the store linearization program transformation. Enabling store linearization resulted in most of the performance degradation observed in earlier tests ($\approx$ 3.5%), with the hardware changes contributing negligible additional overhead ($\approx$ 0.2%). This suggests that performance can be improved further by optimizing the store linearization strategy.
Microbenchmarks.
To assess the impact of our store and load pipeline changes we microbenchmark stores and loads between CHERI-Flute64 in purecap mode and MonCHÉRI-Flute64 in Write-before-Execute + purecap mode. The store microbenchmark writes a 256-element array, recording total ticks. The load microbenchmark, we measured read from an of equal size. Here we report the difference in mean times for the experiment over 10 repetitions. We observed a negligible difference between CHERI-Flute64 and MonCHÉRI-Flute64: $\approx 5$ ticks for the load and $\approx 30$ ticks for the store benchmark.
@@
Finally, we evaluated the impact of Mon CHÉRI on the TLSF allocator through microbenchmarks that allocate and free 1 MB of memory in chunks ranging from 32 bytes to 4 KiB. The results, shown in Figure 5, indicate that the Write-before-Read-enhanced TLSF allocator introduces negligible performance overhead: $\approx$ 0.1% compared to purecap and $\approx$ 1.5% (g.m.) compared to no-cap. The overhead of Write-before-Read + purecap is constant regardless of allocation size, while the overhead of zero-initialization increases linearly with allocation size.
@@
Figure 5: TLSF allocator microbenchmark. Bars show overhead relative to TLSF in MonCHÉRI-Flute64 nocap mode.
6 Related Work
@@
   Uninitialized capabilities [25] *4 ? *3
   Capstone [27] *4 ? $\approx\leavevmode\nobreak\ 50\%$5 *3
   Mon CHÉRI $\approx$ 3.5% /\ /\ / 7%6 *3